40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 7 18:47:36 2011, Jenne, Update, RF System, RAM Mon is now being demodulated 
    Reply  Thu Dec 8 00:04:50 2011, rana, Update, IOO, RAM Mon is now being demodulated 
       Reply  Thu Dec 8 14:47:28 2011, Jenne, Update, IOO, EOM aligned to minimize RAM 
    Reply  Thu Dec 8 00:45:13 2011, Koji, Update, RF System, 4ch demod is ready 
       Reply  Thu Dec 8 22:44:55 2011, Koji, Update, RF System, 4ch demod test result RF_DEMOD_TEST_111208.pdf
Message ID: 6082     Entry time: Wed Dec 7 18:47:36 2011     Reply to this: 6084   6086
Author: Jenne 
Type: Update 
Category: RF System 
Subject: RAM Mon is now being demodulated 

There were 2 open outputs on the splitter in the RAMmon (formerly known as Stochmon) box underneath the BS oplev table.  The input to the splitter comes from the Thorlabs PD that we're using as our RAM monitoring PD.  2 of the outputs go to the RMS detection of 11 and 55 MHz.  Now the other 2 (previously terminated) outputs go over to the LSC rack via SMA cable.  The signal on both channels is ~200mV pk-pk, so -10dBm.  One is plugged into the AS11 demod board (which didn't have a PD input yet), and the other goes to POP55's demod board, so POP55 is not what you think it is for now.

Koji is working on checking out the Rich box, which has 4 demodulators, which we will use eventually.  Right now we're just using the already-plugged-in demod boards so we can start looking at some trends of RAM.  We're going to need to find some channels when we're ready for the switchover.

Zach is nearing completion of the mini-update to the temp sensing system.  Once we have the new more sensitive temp sensor in place, we can have a look-see at the similarities between EOM temperature and RAM levels.

ELOG V3.1.3-